OpenStudy (anonymous):

A street that is 146 m long is covered in snow. City workers are using a snowplow to clear the street. The snowplow has tires that are 1.3 m in diameter. How many times does a tire have to turn in traveling the length of the street? Use the value 3.14 for π . Round your answer to the nearest tenth. Do not round any intermediate steps.

Vocaloid (vocaloid):

number of turns = length of street/circumference of tire

OpenStudy (anonymous):

soooo 146/4.084cm = 0.357492654 meters ?

OpenStudy (anonymous):

or isnt its c=pi*d ?

Vocaloid (vocaloid):

circumference = pi*diameter = (3.14)*1.3 = __ m fill in the blank

OpenStudy (anonymous):

i have 36 for the answer is that correct?

OpenStudy (anonymous):

35.766781 rounded to 36

Vocaloid (vocaloid):

not quite, the question wants you to round to the nearest tenth

OpenStudy (anonymous):

35.8

Vocaloid (vocaloid):

great!

OpenStudy (anonymous):

thank you